# RAMAC: Multimodal Risk-Aware Offline Reinforcement Learning and the Role of Behavior Regularization

### v1

@@ Abstract
In safety-critical domains where online data collection is infeasible, offline reinforcement learning (RL) offers an attractive alternative but only if policies deliver high returns without incurring catastrophic lower-tail risk. Prior work on risk-averse offline RL achieves safety at the cost of value conservatism and restricted policy classes, whereas expressive policies are only used in risk-neutral settings. Here, we address this gap by introducing the Risk-Aware Multimodal Actor-Critic (RAMAC) framework, which couples an expressive generative actor with a distributional critic. The RAMAC differentiates composite objective combining distributional risk and BC loss through the generative path, achieving risk-sensitive learning in complex multimodal scenarios. We instantiate RAMAC with diffusion and flow-matching actors and observe consistent gains in $\mathrm{CVaR}_{0.1}$ while maintaining strong returns on most Stochastic-D4RL tasks. Code: https://github.com/KaiFukazawa/RAMAC.git
@@ 1 Introduction
In high-stakes applications such as autonomous driving, robotics, finance, and healthcare, where real-life explorations may lead to catastrophic consequences, offline RL offers a safe approach for generating policies that not only maximize long-horizon returns but also tightly control risk (Levine et al., 2020). Recent expressive generative policies (Wang et al., 2023; Park et al., 2025; Koirala & Fleming, 2025) can capture multimodal behavior and thus excel in achieving high expected return, yet their primary use has been limited to risk-neutral settings. Conversely, existing risk-averse algorithms ensure safety by enforcing conservatism or restricted policy classes (Kumar et al., 2020; Urpí et al., 2021; Ma et al., 2021). This paper asks: Can we obtain safety without sacrificing expressiveness?
@@
We answer in the affirmative by proposing the Risk-Aware Multimodal Actor-Critic (RAMAC) framework (Fig. 1). RAMAC couples an expressive generative actor with a distributional critic and differentiates a combination of behavioral cloning (BC) and distributional risk (instantiated with Conditional Value-at-Risk (CVaR)) gradients through the generative process (Di Castro et al., 2012; Chow et al., 2015), thereby unifying high expressiveness with robust tail-risk control, and reducing the out-of-distribution (OOD) action.
@@
Prior offline-RL approaches can be organized by mechanism: (i) Policy regularization constrains the policy to the data manifold via divergence minimization or policy priors, improving stability but often sacrificing policy expressiveness on complex tasks with risk-neutral examples such as (Fujimoto et al., 2019; Wu et al., 2019; Kumar et al., 2019; Fujimoto & Gu, 2021) and risk-aware methods with prior-anchored perturbation designs such as (Urpí et al., 2021; Chen et al., 2024). (ii) Value conservatism reduces optimistic extrapolation, but can underestimate the value of infrequent yet high-return in-distribution modes due to global pessimism and data imbalance in both risk-neutral ((Kumar et al., 2020)) and risk-aware instances ((Ma et al., 2021)). (iii) Model-based pessimism bounds transition uncertainty with ensembles and penalties, at the cost of compounding model errors at scale again under both risk-neutral ((Yu et al., 2020; 2021; Rigter et al., 2022)) and risk-aware ((Rigter et al., 2023)) settings. (iv) Expressive generative policies faithfully clone multimodal behavior and achieve state-of-the-art mean returns, but limited use only in risk-neutral applications (Chen et al., 2021; Janner et al., 2022; Ajay et al., 2022; Wang et al., 2023; Hansen-Estruch et al., 2023; Park et al., 2025) including closest concurrent works pairing diffusion with distributional critics (Anonymous, 2025; Liu et al., 2025).
Despite compelling results from expressive models (e.g., diffusion, flow matching) in risk-neutral RL, their potential in offline risk-aware RL remains largely untapped.
@@
Here, we aim to leverage the advantages of expressive policies without compromising risk-aversion or increasing the OOD action rate. To this end, inspired by the success of risk-neutral expressive policies such as  (Wang et al., 2023), RAMAC optimizes a joint objective composed of a BC and CVaR. The direct empirical BC term reduces BC estimation error effect observed in methods such as  (Nair et al., 2020) and hence, reduces the OOD visitation rate which is a critical issue in offline RL. The CVaR term in the objective removes risk-blindness (Fig. 2). We show that RAMAC yields high expected return while minimizing risk on complex multimodal offline benchmarks.
@@
  • Risk-aware expressive policy learning: We leverage expressive policies in the context of risk-aware RL and present two instantiations: RADAC (diffusion) and RAFMAC (flow matching).
  • Theoretical insight: We provide a theoretical discussion on how a BC regularized objective can improve performance in the context of offline RL. Driven by our theoretical results, we conjecture on one possible mechanism through which expressive policies can further improve performance.
  • Experimental evaluation: On the Stochastic-D4RL, our two instantiations (RADAC / RAFMAC) outperform baselines on CVaR while maintaining competitive mean return on most tasks.
@@ 2 Preliminaries
We consider a finite-horizon Markov Decision Process (MDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma,H)$ (Sutton et al., 1998) and a fixed offline dataset $\mathcal{D}=\{(s_{t},a_{t},r_{t},s_{t+1})\}_{i=1}^{N}$ collected by an unknown behavior policy $\beta$ (Prudencio et al., 2023). Let $\operatorname{supp}(\mathcal{D})$ denote the dataset’s empirical state–action support. The objective is to learn a policy $\pi$ that maximizes the expected return $J(\pi)=\mathbb{E}_{\pi,P}[\sum_{t=0}^{H-1}\gamma^{t}r_{t}]$ without extra environment interaction. The central challenge is distributional shift (i.e., OOD): When $\pi$ visits $(s,a)\notin\operatorname{supp}(\mathcal{D})$, value estimates extrapolate and can be unsafe (Kumar et al., 2020). Prior work alleviates this issue with behavior regularization, conservative critics, or model-based pessimism.
@@
A large family of offline methods uses an actor–critic with an explicit proximity term to the behavior policy (Nair et al., 2020; Fujimoto & Gu, 2021; Wu et al., 2019; Kumar et al., 2019):
@@
We expand this variant into the behavior-regularized distributional actor-critic framework in which value term optimizes a coherent risk measure in place of the mean Q, yielding risk-aware updates while retaining the same BC regularize.
@@
To move beyond expectation, distributional RL learns the return law via the distributional Bellman operator (Bellemare et al., 2017)
@@
and often parameterizes the inverse CDF $Z_{\phi}(s,a;\tau)$ with an Implicit Quantile Network (IQN) (Dabney et al., 2018). Access to quantiles enables coherent risk measure $\mathcal{D}(\cdot)$, such as the CVaR. For a risk level $\alpha\in(0,1]$, the CVaR admits a dual form (Rockafellar et al., 2000):
$\mathrm{CVaR}_{\alpha}(X)\;=\;\inf_{\begin{subarray}{c}q\ll P,\,0\leq\frac{dq}{dP}\leq\frac{1}{\alpha}\end{subarray}}\mathbb{E}_{q}[X].$ (4)
and the integral form (used for actor gradients) is $\mathrm{CVaR}_{\alpha}(X)=\frac{1}{\alpha}\int_{0}^{\alpha}F_{X}^{-1}(\tau)\,d\tau$.
@@
Recent actors generate an action by evolving a latent $z\!\sim\!\mathcal{N}(0,I)$ along a differentiable path (Janner et al., 2022; Wang et al., 2023). We focus on the two families:
(i) Diffusion policies follow a reverse-time SDE (Song et al., 2021b),
@@
and (ii) Flow-matching policies solve a deterministic ODE (Lipman et al., 2023),
@@
The entire map $\psi_{\theta}:s,z\mapsto a$ is differentiable, enabling guidance from a critic through denoising/integration steps. Prior work typically injects expected-value signals; our framework instead inject distributional risk signals.
@@ 3 Risk-Aware Multimodal Actor-Critic (RAMAC)
We now introduce the Risk-Aware Multimodal Actor-Critic (RAMAC). At its core, RAMAC operates in two stages: First, a distributional critic parameterized as an IQN learns the full conditional distribution of returns. Second, a generative actor, instantiated as either a diffusion policy or a flow-matching policy, is guided jointly by two terms in the objective function: (i) BC term that constrains the policy to the data manifold and (ii) CVaR term extracted from the critic’s lower tail. The latter pushes the probability mass away from low-probability, catastrophic regions and preserves the high-reward modes (Fig. 2). The former acts as a regularizer and limits OOD visitation rate.
@@ 3.1 Distributional Critic
Risk-sensitive objectives such as CVaR require access to the entire return distribution. We therefore adopt a distributional critic $Z_{\phi}$ via IQN (Dabney et al., 2018), building on the distributional Bellman operator in Eq. 3 Bellemare et al. (2017). We minimize a distributional Bellman residual with a quantile-Huber loss (with $\kappa\!=\!1$):
@@ 3.3 Behavior-regularized Objective
The complete policy objective balances risk aversion with fidelity to the offline dataset. It combines the risk term with a standard behavior cloning (BC) loss, $\mathcal{L}_{\text{BC}}(\theta)$:
@@
where $\eta$ is a hyperparameter. In this work, we demonstrate both diffusion (RADAC) and flow (RAFMAC) variants. We show a pseudocode for RAMAC in Algorithm 1 and describe the full implementation details in App. C
@@
4 Behavior Regularization
Prior work has demonstrated the importance of behavior regularization in offline RL due to its ability to constrain the learned policy to the data manifold, curb value extrapolation, and stabilize improvement under distributional shift. A commonly adopted regularization scheme in offline risk-aware RL is prior-anchored perturbation method (e.g. ORAAC, UDAC) [^2] For simplicity and consistency with our experiments, we will refer to UDAC as ORAAC–Diffusion (Urpí et al., 2021; Chen et al., 2024), which uses a linear mixing of a pretrained BC policy with the RL actor (perturbation). Here, we first discuss the limitation of this regularization approach. We then demonstrate the advantages of our adopted scheme, namely, behavior-regularized objective method (shown in Eq. 11).

### v2

@@ Abstract
In safety-critical domains where online data collection is infeasible, offline reinforcement learning (RL) offers an attractive alternative but only if policies deliver high returns without incurring catastrophic lower-tail risk. Prior work on risk-averse offline RL achieves safety at the cost of value or model-based pessimism, and restricted policy classes that limit policy expressiveness, whereas diffusion/flow-based expressive generative policies trained with a behavioral-cloning (BC) objective have been used only in risk-neutral settings. Here, we address this gap by introducing the Risk-Aware Multimodal Actor-Critic (RAMAC), which couples an expressive generative actor with a distributional critic and, to our knowledge, is the first model-free approach that learns risk-aware expressive generative policies. RAMAC differentiates a composite objective that adds a Conditional Value-at-Risk (CVaR) term to a BC loss, achieving risk-sensitive learning in complex multimodal scenarios. Since out-of-distribution (OOD) actions are a major driver of catastrophic failures in offline RL, we further analyze OOD behavior under prior-anchored perturbation schemes from recent BC-regularized risk-averse offline RL. This clarifies why a behavior-regularized objective that directly constrains the expressive generative policy to the dataset support provides an effective, risk-agnostic mechanism for suppressing OOD actions in modern expressive policies. We instantiate RAMAC with a diffusion-based actor, using it both to illustrate the analysis in a 2-D risky bandit and to deploy OOD-action detectors on Stochastic-D4RL benchmarks, empirically validating our insights. Across these tasks, we observe consistent gains in $\mathrm{CVaR}_{0.1}$ while maintaining strong returns. Our implementation is available at GitHub: https://github.com/KaiFukazawa/RAMAC.git
@@ 1 Introduction
In high-stakes applications such as autonomous driving, robotics, finance, and healthcare, where real-world explorations may lead to catastrophic consequences, offline RL offers a safe approach for generating policies that not only maximize long-horizon returns but also tightly control risk [levine2020offline]. Recent expressive generative policies [koirala2025flow; park2025flow; wang2022diffusion] can capture multimodal behavior and thus excel in achieving high expected return, yet their primary use has been limited to risk-neutral settings. Conversely, existing risk-averse algorithms ensure safety by enforcing conservatism or restricted policy classes, which constrain policy expressiveness [kumar2020conservative; ma2021conservative; urpi2021risk]. This paper asks: Can we obtain safety without sacrificing expressiveness?
@@
We answer in the affirmative by proposing the Risk-Aware Multimodal Actor-Critic (RAMAC) framework (Fig. 1). RAMAC couples an expressive generative actor with a distributional critic and differentiates a combination of BC and distributional risk (instantiated with CVaR) gradients through the generative process [chow2015risk; di2012policy]. This unifies high expressiveness with robust tail-risk control while directly constraining the generative policy to the data support, addressing two central safety concerns in offline RL: catastrophic tail outcomes and out-of-distribution (OOD) actions.
@@
Prior offline-RL approaches can be organized by mechanism: (i) Policy regularization constrains the policy to the data manifold via divergence minimization or policy priors, improving stability but often sacrificing policy expressiveness on complex tasks with risk-neutral examples such as [fujimoto2021minimalist; fujimoto2019off; kumar2019stabilizing; wu2019behavior] and risk-aware methods with prior-anchored perturbation designs such as [chen2024uncertainty; urpi2021risk]. (ii) Value conservatism reduces optimistic extrapolation, but can underestimate the value of infrequent yet high-return in-distribution modes due to global pessimism and data imbalance in both risk-neutral [kumar2020conservative] and risk-aware instances [ma2021conservative]. (iii) Model-based pessimism bounds transition uncertainty with ensembles and penalties, at the cost of compounding model errors at scale again under both risk-neutral [rigter2022rambo; yu2020mopo; yu2021combo] and risk-aware [rigter2023one] settings. (iv) Expressive generative policies faithfully clone multimodal behavior and achieve state-of-the-art mean returns, but limited use only in risk-neutral applications [hansen2023idql; kang2023efficient; koirala2025flow; park2025flow; wang2022diffusion] including closely related concurrent works that pair diffusion with distributional critics [liu2025distributional; zhang2025d2]. In these methods, expressive diffusion/flow policies trained with BC-style objectives have been widely adopted in offline RL.
Despite compelling results from these expressive generative policies in risk-neutral RL, their potential in offline risk-aware RL remains largely untapped. Among behavior-regularized risk-averse methods, the most prominent approaches that leverage expressive priors rely on prior-anchored perturbation [chen2024uncertainty; urpi2021risk], which trade away much of the multimodal capacity and, as we show theoretically, can still incur OOD actions.
@@
Here, we aim to leverage the advantages of expressive policies without compromising risk-aversion or increasing the OOD action rate. To this end, inspired by the success of risk-neutral expressive policies such as [wang2022diffusion], RAMAC optimizes a joint objective composed of a BC loss on the generative actor and a CVaR objective on the return distribution. Unlike prior diffusion/flow-based offline RL methods, which rely on BC-style objectives primarily as an empirical safeguard against extrapolation [chen2024diffusion], our BC term induces a forward-KL upper bound on the probability of taking off-support actions, directly linking behavior regularization and tail-risk control (Sec. 4). Combined with the CVaR-based risk term, this yields a principled behavior-regularized objective for expressive risk-aware policies. To our knowledge, this is the first objective-level characterization of how a BC-regularized expressive policy controls OOD behavior. We empirically show that RAMAC yields high expected return while minimizing risk on complex multimodal offline benchmarks.
@@
  • Risk-aware expressive policy learning: We leverage expressive policies in the context of risk-aware RL by introducing RAMAC, a model-free framework for learning risk-aware expressive generative policies. Our primary instantiation is a diffusion-based actor, RADAC (Risk-Aware Diffusion Actor-Critic).
  • Theoretical insight:
    We provide a geometric analysis of OOD behavior in offline RL, showing that prior-anchored perturbation schemes can still place probability mass on off-support actions even with expressive priors. In contrast, for expressive generative policies, we derive a forward-KL upper bound on the per-state OOD action probability, explaining why a behavior-regularized objective on the generative actor is an effective, risk-agnostic way to control OOD visitation.
  • Experimental evaluation: On Stochastic-D4RL benchmarks, RADAC outperforms baselines on CVaR while maintaining competitive mean return on most tasks. We additionally report a flow-matching variant, RAFMAC (Risk-Aware Flow-Matching Actor-Critic), in App. E.2, which shows similar trends. We also (i) visualize our geometric analysis in a 2-D risky bandit that contrasts risk-aware prior-anchored perturbation with expressive priors and risk-aware expressive generative policies, and (ii) use OOD-action detectors on Stochastic-D4RL to empirically validate the theoretical insights.
@@ 2 Preliminaries
We consider a finite-horizon Markov Decision Process (MDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma,H)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition kernel $P(\cdot\mid s,a)$, reward function $r(s,a)$, discount factor $\gamma\in(0,1)$, and horizon $H\in\mathbb{N}$ [sutton1998reinforcement]. In offline RL, the learner is given only a static dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i},s^{\prime}_{i})\}_{i=1}^{N}$ collected by some unknown behavior policy $\beta$, and cannot further interact with the environment [prudencio2023survey]. Let $\operatorname{supp}(\mathcal{D})$ denote the dataset’s empirical state–action support. The objective is to learn a policy $\pi$ that maximizes the expected return $J(\pi)=\mathbb{E}_{\pi,P}[\sum_{t=0}^{H-1}\gamma^{t}r_{t}]$ without extra environment interaction. The central challenge is distributional shift (i.e., OOD): When $\pi$ visits $(s,a)\notin\operatorname{supp}(\mathcal{D})$, value estimates $Q(s,a)$ extrapolate and can become arbitrarily inaccurate [kumar2020conservative]. Policies that place non-negligible mass on such OOD actions may therefore suffer catastrophic failures at deployment. Prior work alleviates this issue with behavior regularization, conservative critics, or model-based pessimism.
@@
A large family of offline methods uses an actor–critic with an explicit proximity term to the behavior policy [fujimoto2021minimalist; kumar2019stabilizing; nair2020awac; wu2019behavior]. A representative actor–critic objective takes the form:
@@
Here the second term $\alpha\log\pi_{\theta}(a\mid s)$ plays the role of a behavior regularizer: it is typically instantiated as a KL- or behavioral-cloning proximity term that keeps $\pi_{\theta}(\cdot\mid s)$ close to the behavior policy $\beta(\cdot\mid s)$. Empirically, BRAC-style objectives have turned out to be surprisingly strong in offline RL  [tarasov2023revisiting]. In this work, we extend this behavior-regularized pattern to a distributional actor–critic in which the critic is expanded into a distributional critic that evaluates a coherent risk measure (CVaR) in place of the mean $Q$.
@@
Standard actor–critic methods including BRAC optimize the expected return by learning the mean action-value function $Q^{\pi}(s,a)=\mathbb{E}[Z^{\pi}(s,a)]$ as shown in Eq. 2. Distributional RL instead models the entire return distribution $Z^{\pi}(s,a)$ [bellemare2017distributional]. The distributional Bellman operator is:
@@
A common parameterization uses an Implicit Quantile Network (IQN) [dabney2018implicit] to approximate the inverse cumulative distribution function (CDF) $Z_{\phi}(s,a;\tau)\approx F^{-1}_{Z^{\pi}(s,a)}(\tau)$ for quantile levels $\tau\in(0,1)$. Access to quantiles enables coherent risk measures $\mathcal{D}(\cdot)$ that emphasize different parts of the return distribution. In this work we focus on the Conditional Value-at-Risk (CVaR), a widely used instantiation of $\mathcal{D}(\cdot)$ as risk-averse objective. For a risk level $\alpha\in(0,1]$, the CVaR admits the integral form used for actor gradients:
$\mathrm{CVaR}_{\alpha}(X)=\frac{1}{\alpha}\int_{0}^{\alpha}F_{X}^{-1}(\tau)\,d\tau.$ (4)
Optimizing $\mathrm{CVaR}_{\alpha}$ encourages policies that trade some mean performance for improved behavior in the worst $\alpha$-fraction of trajectories, which is crucial in safety-critical settings. In RAMAC, the distributional critic provides quantile estimates from which CVaR and its gradients with respect to actions can be computed and backpropagated through the policy.
@@
Recent offline RL methods stay within the behavior-regularized actor-critic template of Eqs. 1 and 2, but replace the simple parametric actor with an expressive conditional generative model  [hansen2023idql; kang2023efficient; koirala2025flow; park2025flow; wang2022diffusion]. Given a state $s$ and latent $z\!\sim\!\mathcal{N}(0,I)$, the policy generates an action $a=\psi_{\theta}(s,z)$ along a differentiable path  [hansen2023idql; wang2022diffusion], while an explicit behavior term keeps $\psi_{\theta}$ close to the dataset action signals. We focus on the two families:
(i) Diffusion policies model a reverse-time stochastic differential equation (SDE) over actions [song2020score],
@@
where a forward noising process gradually corrupts dataset actions into near-Gaussian noise, and the network $f_{\theta}$ learns to reverse this process conditioned on the state $s$.
(ii) Flow-matching policies solve a deterministic ODE [lipman2022flow],
@@
where a neural vector field $v_{\theta}$ transports samples from a simple base distribution to the data distribution along a continuous trajectory. Integrating Eq. 6 from an initial noise sample yields an action conditioned on $s$.
In both cases, the overall map $\psi_{\theta}:(s,z)\mapsto a$ is fully differentiable, and the behavior term encourages $\psi_{\theta}$ to approximate the behavior action distribution itself; the critic then fine-tunes this expressive behavior model using scalar signals. Prior work typically uses expected-value or advantage-based signals from a mean-value critic to update the generative policy (as in Eq. 1), yielding a risk-neutral generative BRAC method. In our case, a distributional critic instead provides explicit distributional risk signals, namely tail-sensitive values such as $\mathrm{CVaR}_{\alpha}(Z^{\pi}(s,a))$ derived from the return distribution (as in Eq. 4), that shape the diffusion/flow policy under the same behavior-regularized template; the exact loss is introduced in Sec. 3.
@@ 3 Risk-Aware Multimodal Actor-Critic (RAMAC)
We now introduce the Risk-Aware Multimodal Actor-Critic (RAMAC). At its core, RAMAC operates in two stages. First, a distributional critic parameterized as an IQN learns the full conditional distribution of returns. Second, a generative actor, instantiated as either a diffusion policy or a flow-matching policy, is guided jointly by two terms in the objective function: (i) BC term that constrains the policy to the data manifold and (ii) CVaR term extracted from the critic’s lower tail. The latter pushes the probability mass away from low-probability, catastrophic regions and preserves the high-reward modes (Fig. 2). The former acts as a regularizer: by keeping $\pi_{\theta}(\cdot\mid s)$ close to the empirical behavior distribution, it limits OOD visitation, mirroring the BRAC-style regularizers in Eq. 1 and connecting directly to our forward-KL OOD bound in Sec. 4.
@@ 3.1 Distributional Critic
Risk-sensitive objectives such as CVaR require access to the entire return distribution. We therefore adopt a distributional critic $Z_{\phi}$ via IQN [dabney2018implicit], building on the distributional Bellman operator in Eq. 3. We minimize a distributional Bellman residual with a quantile-Huber loss (with $\kappa\!=\!1$):
@@ 3.3 Behavior-regularized Objective
The complete policy objective balances risk aversion with fidelity to the offline dataset. We instantiate fidelity through a standard behavior cloning (BC) term that encourages the policy to reproduce the behavior distribution. We define:
$\mathcal{L}_{\text{BC}}(\theta)=-\,\mathbb{E}_{(s,a)\sim\mathcal{D}}\big[\log\pi_{\theta}(a\mid s)\big],$ (11)
which matches the BRAC-style behavior regularizer in Eq. 1 up to a scaling of the coefficient. [^2] For diffusion/flow actors, we implement $\mathcal{L}_{\text{BC}}$ via standard denoising / flow-matching losses. It combines the risk term with a standard behavior cloning (BC) loss, $\mathcal{L}_{\text{BC}}(\theta)$:
@@
where $\eta$ is a hyperparameter. Our primary instantiation is a diffusion policy (RADAC), while an additional flow-matching variant (RAFMAC) is reported in the App. E.2. We show pseudocode for RAMAC in Algorithm 1 and describe the full implementation details in App. D
@@
4 Behavior Regularization in Offline RL
Prior work has demonstrated the importance of behavior regularization in offline RL due to its ability to constrain the learned policy to the data manifold, curb value extrapolation, and stabilize improvement [tarasov2023revisiting]. A commonly adopted regularization scheme in offline risk-aware RL is the prior-anchored perturbation method (e.g. ORAAC, UDAC) [^3] For simplicity and consistency with our experiments, we will refer to UDAC as ORAAC–Diffusion [chen2024uncertainty; urpi2021risk], which uses a linear mixing of a pretrained BC policy with the RL actor (perturbation). Here, we first discuss the limitation of this regularization approach. We then highlight the advantages of our scheme, namely, the behavior-regularized objective in Eq. 12.